# SuperRL: Reinforcement Learning with Supervision to Boost Language Model Reasoning

### v1

@@ Abstract
Large language models (LLMs) are increasingly applied to complex reasoning tasks, where rich offline data—such as expert-annotated solutions or distilled traces—are often available. However, conventional reinforcement learning (RL) fails to exploit this data effectively, especially under sparse-reward environment. We propose SuperRL, a unified training framework to solve the problem. SuperRL employs an Adaptive Switch to detect reward sparsity and invokes a Hybrid Actor when rewards are sparse, blending policy gradients with offline supervision to stabilize learning. Experiments across diverse reasoning benchmarks demonstrate that SuperRL consistently outperforms standard RL, delivering improved sample efficiency, generalization, and robustness under sparse rewards. Refer to caption Figure 1: Overview of SuperRL. To adaptively optimize under varying reward conditions, SuperRL introduces an Adaptive Switch that conducts a brief rollout probe at the start of training to determine whether the environment provides dense or sparse rewards. In dense-reward settings, the model applies standard reinforcement learning (RL) updates to take full advantage of abundant feedback. When rewards are sparse, control is handed to a Hybrid Actor that augments the policy-gradient objective with a supervised fine-tuning (SFT) loss derived from high-quality offline trajectories. This hybrid update stabilizes learning and improves sample efficiency by compensating for insufficient or delayed reward signals. By dynamically switching objectives based on reward density, SuperRL effectively combines exploration with reliable offline supervision to support robust reasoning.
@@ 1 Introduction
Despite recent progress, online reinforcement learning methods such as PPO and GRPO still face limitations in reasoning tasks. [8, 10] First, these inherently on-policy methods update exclusively from trajectories generated by the current policy and therefore struggle to leverage high-quality offline data—such as human-annotated or distilled reasoning traces—that encode valuable prior knowledge but lie outside the policy distribution. [11, 12] Second, in sparse-reward environments, online rollouts seldom produce enough successful trajectories to supply reliable learning signals, making it difficult to bootstrap coherent reasoning. [13, 14] Ultimately, because offline data provide numerous successful trajectories with clear guidance while online rollouts frequently fail and offer little feedback, reinforcement learning struggles to acquire robust reasoning capabilities in sparse-reward environments.
@@
To enhance language model reasoning across both dense and sparse reward regimes, we propose SuperRL, a unified training framework built around two key components: the Adaptive Switch and the Hybrid Actor. The Adaptive Switch begins training with a brief rollout probe to determine whether the environment provides dense or sparse rewards. In dense-reward environment, SuperRL performs standard reinforcement learning updates to take advantage of abundant feedback. When rewards are sparse, control is handed over to the Hybrid Actor, which integrates reinforcement learning with supervised fine-tuning on high-quality offline reasoning traces—thus enabling the model to benefit from reliable “correct thinking” examples that standard on-policy methods are unable to effectively utilize. SuperRL is able to maintain robust, generalizable reasoning performance under sparse-reward conditions where traditional RL struggles to learn from valuable offline data.
As shown in Table 1, on sparse-reward datasets, SuperRL outperforms vanilla RL by large margins, +6.9% on LIMO, +15.0% on OpenR1, and +64.2% on HiTab. It also achieves notable generalization improvements on out-of-distribution datasets, demonstrating strong robustness under sparse-reward conditions. 4.3
Our main contributions are:
  • We propose SuperRL, a unified training framework that adaptively switches between Vanilla and Hybrid RL actors based on the density of reward signals, enabling language models to reason effectively under both dense and sparse reward settings.
  • We design a Hybrid Actor that interleaves policy-gradient updates with supervised fine-tuning on high-quality offline reasoning traces, striking a balance between exploration and reliable learning signals when rewards are sparse.
  • We conduct comprehensive experiments across diverse datasets, model scales, and training regimes, demonstrating the robustness, effectiveness, and generalizability of the SuperRL framework in enhancing reasoning performance.
@@ 2.2 Reinforcement Learning for LLMs
Reinforcement learning from human feedback [24]—as popularized by InstructGPT—uses on-policy methods like PPO to fine-tune language models against a learned reward signal, but doing so requires costly fresh sampling, a separate value network and careful KL-penalty tuning, and remains sample-inefficient when rewards are sparse. [10, 25, 22] GRPO simplifies PPO by ditching the critic and computing advantages via within-prompt reward normalization across multiple candidates, yet it still depends on large volumes of model-generated outputs and can wander from the pre-trained behavior in low-reward regimes. These limitations—computational overhead, sample inefficiency and instability under sparse feedback—have spurred interest in leaner or hybrid approaches that retain the simplicity of policy gradients while injecting external guidance to stabilize and steer training.[8, 22]
@@ 2.3 Unifying SFT and RL: Towards Hybrid Optimization
Unlike prior methods that apply supervised fine-tuning (SFT) and reinforcement learning (RL) in separate stages, SuperRL unifies them by injecting SFT signals directly into the RL loss function. This dynamic integration enables the model to simultaneously leverage the strengths of high-quality offline demonstrations and the exploratory benefits of online rollouts. By guiding policy updates with both supervised and reinforcement signals, SuperRL significantly improves reasoning capabilities and generalization performance. This unified optimization paradigm offers a principled framework for incorporating offline knowledge into reinforcement learning.
@@ 3 Methodology
We present SuperRL, a unified training framework comprising two key components—Adaptive Switch and Hybrid Actor—designed to adaptively train language models across both dense and sparse reward environments. Unlike conventional SFT-RL pipelines, SuperRL dynamically monitors reward density and selects the appropriate training strategy: standard on-policy updates for dense rewards, or an integrated RL+SFT approach for sparse rewards. This preserves the sample efficiency of offline data while retaining the exploration benefits of reinforcement learning.
The Adaptive Switch conducts an initial rollout probe to assess reward density. If the average reward exceeds a predefined threshold, the model proceeds with vanilla actor–critic training. Otherwise, control is passed to the Hybrid Actor. This dynamic mechanism eliminates the need for manual scheduling and ensures automatic adaptation to varying reward structures, promoting stability and efficiency.
When activated, the Hybrid Actor blends policy-gradient updates with supervised fine-tuning on high-quality offline reasoning traces. By combining RL and SFT losses, the Hybrid Actor enhances learning in sparse-reward settings, improving convergence and robustness.
Together, these components form a single end-to-end training loop, unifying pre-training and fine-tuning. Under dense rewards, SuperRL operates as standard RL; under sparse rewards, it augments training with offline demonstrations. Full implementation and hyperparameter details are provided in this section.
3.1 Discussion of Adaptive Fusion of SFT and RL
We propose a adaptive fusion approach that integrates SFT and RL signals within each update step, dynamically weighted by uncertainty. Unlike traditional sequential paradigms where models undergo SFT followed by RL, our method continuously blends both objectives throughout training. This design captures the best of both worlds: the strong behavioral alignment of SFT and the exploratory capacity of RL. SFT provides rapid convergence toward expert-like behavior but often overfits and lacks robustness in unseen scenarios, especially under sparse rewards. [15, 16, 17] RL, on the other hand, promotes exploration and generalization but suffers from high variance and instability in low-signal environments. [17] Sequential SFT-RL pipelines often experience severe misalignment due to abrupt objective shifts, sometimes leading to catastrophic forgetting that RL alone cannot rectify. By contrast, our soft fusion strategy avoids such abrupt transitions. [19, 20, 21] Through uncertainty-based adaptive weighting, each update jointly incorporates both demonstration-driven and reward-driven signals, enabling more stable and efficient learning, particularly in sparse-reward settings.
The design yields two key benefits:
  • Stable early-stage learning: In the presence of high uncertainty, the model leans more heavily on SFT signals, which encapsulate rich prior knowledge. This mitigates premature exploration and stabilizes initial training dynamics.
  • Resilient objective adaptation: The framework continuously balances between SFT and reinforcement learning, dynamically prioritizing the more informative signal. This soft objective selection enables smooth adaptation to evolving task requirements, avoiding brittle transitions or model collapse.
We posit that this dynamic and continuous balance—rather than discrete handovers—forms the foundation of the model’s robustness and strong empirical performance in sparse-reward, reasoning-intensive tasks (see Appendix E for a comparative analysis of soft versus hard switching strategies).
3.2 Uncertainty-Weighted Hybrid Actor Design
To unify supervised and reinforcement learning objectives within a single optimization loop, we introduce an uncertainty-weighted hybrid actor. Concretely, we introduce two log-variance parameters, $\sigma_{\mathrm{pg}}$ and $\sigma_{\mathrm{sft}}$, which respectively modulate the contributions of the policy-gradient loss $L_{\mathrm{actor}}$ and the SFT loss
$L_{\mathrm{sft}}(\theta)=\mathbb{E}_{(x,y)\sim D_{\mathrm{sft}}}\bigl{[}\mathrm{CrossEntropy}\bigl{(}\pi_{\theta}(y\mid x),y\bigr{)}\bigr{]}.$
Our hybrid objective is then written as
$L_{\mathrm{Hybrid}}=\exp(-2\sigma_{\mathrm{pg}})\,L_{\mathrm{actor}}\;+\;\exp(-2\sigma_{\mathrm{sft}})\,L_{\mathrm{sft}}\;+\;\sigma_{\mathrm{pg}}+\sigma_{\mathrm{sft}},$
where $L_{\mathrm{actor}}$ can be instantiated by any policy-gradient method—e.g. PPO with its clipped surrogate
$L_{\mathrm{ppo}}=\mathbb{E}_{t}\bigl{[}\min(r_{t}A_{t},\;\mathrm{clip}(r_{t},1-\epsilon,1+\epsilon)A_{t})\bigr{]}-\beta_{\mathrm{ent}}\,\mathbb{E}_{t}[\mathcal{H}(\pi_{\theta})]$
or GRPO, which further normalizes advantages within reasoning-template groups before clipping.
During training, we mix supervised fine-tuning (SFT) and reinforcement learning (RL) signals within each backward pass by interleaving high-quality demonstration data with on-policy rollouts. This design allows the model to learn from both expert behavior and exploration at the same time, ensuring that the two learning signals are tightly coupled throughout optimization.
To balance these signals, we use two learnable uncertainty parameters, $\sigma_{\mathrm{pg}}$ and $\sigma_{\mathrm{sft}}$, which control how much each loss contributes. Noisy or unstable signals are automatically down-weighted through the exponential terms $\exp(-2\sigma)$, while the additive $\sigma$ terms prevent the model from becoming overconfident. Because these weights are learned dynamically, the model adapts naturally to differences in signal quality without needing explicit switching rules.
This uncertainty-based fusion leads to more stable training, faster convergence, and stronger generalization—especially in sparse-reward, reasoning-heavy environments. By grounding updates in high-quality offline data while still allowing for on-policy exploration, our hybrid actor achieves better performance with a simpler and more unified training process.
3.3 Adaptive Switch Design
To dynamically select the most appropriate training strategy during the early stages of learning, we design a modular decision mechanism called the Adaptive Switch. This component allows SuperRL to automatically tailor its optimization behavior based on the reward characteristics of the environment, thereby avoiding rigid, one-size-fits-all solutions and improving generalization across diverse tasks.
Reward Density Assessment. Before formal training begins, SuperRL performs a lightweight rollout probe to assess the reward structure of the task. During this initial phase, the system executes a small number of pure RL training steps and monitors two key indicators: the average reward observed at each step and the number of times this average reward increases over the rollout window. Additionally, it calculates the average of the most recent few reward values to evaluate short-term learning progress. If both the increase count and the recent average reward fall below predefined thresholds, the task is classified as having sparse rewards; otherwise, it is treated as dense reward.
Based on this early analysis, the Adaptive Switch chooses the most suitable training actor. For dense-reward tasks, it selects a standard reinforcement learning actor, which benefits from the efficiency and alignment of pure policy-gradient updates. For sparse-reward tasks, it activates the proposed uncertainty-weighted hybrid actor, which incorporates supervised fine-tuning signals into every update.  4.2  4.3
We empirically determine effective threshold values for different batch size settings. When the batch size is larger than 32, we use 10 rollout steps, track the last 10 steps for recent average reward, and set the thresholds to 3 reward increases and a recent average reward of 0.1. For smaller batch sizes, we increase the number of rollout steps to 50, keep the recent average over 10 steps, and raise the thresholds to 20 increases and a recent average reward of 0.2. These values were selected based on extensive experiments across benchmarks such as GSM8K, MetaMath, LIMO, and so on. 4.5
The modular nature of the Adaptive Switch makes it easy to interpret and customize. Practitioners can examine reward signals and actor decisions directly or override them using domain knowledge. This design not only streamlines system tuning but also enhances transparency and control.
Further extensions and discussions on the Adaptive Switch are provided in Appendix A.

### v2

@@ Abstract
Large language models (LLMs) are increasingly applied to complex reasoning tasks, where rich high-quality offline data—such as expert-annotated solutions—are often available. However, standard reinforcement learning (RL) struggles in sparse-reward settings and fails to fully leverage offline supervision. We propose SuperRL, a unified training framework that adaptively switches between RL and supervised fine-tuning (SFT) . For any data instance where all sampled trajectories yield zero reward, indicating a lack of gradient signal, SuperRL triggers a fallback to supervised fine-tuning using high-quality offline data. Experiments across diverse reasoning benchmarks demonstrate that SuperRL outperforms standard RL, delivering improved sample efficiency, generalization, and robustness. Refer to caption Figure 1: Overview of SuperRL. SuperRL is a unified training framework that adaptively combines RL and SFT based on reward signal. During training, for each input, the model samples multiple rollouts and computes their rewards. If at least one trajectory receives a nonzero reward, standard RL updates are applied using policy gradients. If all trajectories yield zero reward—indicating an absence of learning signal—SuperRL falls back to supervised fine-tuning using high-quality demonstrations.
@@ 1 Introduction
Despite recent progress, online reinforcement learning methods such as PPO and GRPO still face limitations in reasoning tasks. [8, 10] First, these inherently on-policy methods update exclusively from trajectories generated by the current policy and therefore struggle to leverage high-quality offline data—such as human-annotated or distilled reasoning traces—that encode valuable prior knowledge but lie outside the policy distribution. [11, 12] Second, in sparse-reward environments, online rollouts seldom produce enough successful trajectories to supply reliable learning signals, making it difficult to bootstrap coherent reasoning. [13, 14] Ultimately, because offline data provide numerous successful trajectories with clear guidance while online rollouts frequently fail and offer little feedback, reinforcement learning struggles to acquire robust reasoning capabilities.
@@
To enhance language model reasoning across both dense and sparse reward regimes, we propose SuperRL, a unified training framework that dynamically switches between RL and SFT at the instance level. For each training input, the model samples multiple rollouts and evaluates their rewards. If any rollout receives a nonzero reward, policy gradient updates are applied. Otherwise—when all rollouts fail—SuperRL falls back to SFT on expert demonstrations. This simple yet effective fallback mechanism allows the model to benefit from high-quality offline data whenever online learning provides insufficient signal. While online rollouts frequently fail to offer informative supervision, offline datasets contain abundant successful trajectories that are often underutilized. By adaptively leveraging these trajectories, SuperRL closes this gap and facilitates more stable and effective learning. SuperRL is entirely data-driven and achieves strong generalization across reasoning tasks.
Our main contributions are as follows:
  • SFT-guided Reinforcement Learning. We propose a simple yet powerful mechanism to integrate SFT signals into the RL process. This hybrid approach enables the model to fall back on high-quality offline supervision in situations where reward signals are sparse, delayed, or entirely absent, thereby significantly enhancing the learning signal in complex reasoning environments.
  • The SuperRL Framework. We introduce SuperRL, a training framework that dynamically switches between RL and SFT at the instance level. The switching is guided by reward feedback, allowing the system to adaptively determine the most appropriate optimization strategy for each input.
  • Comprehensive Empirical Validation. We conduct extensive experiments across multiple reasoning benchmarks and model scales. Results show that SuperRL outperforms standard RL baselines in most settings, yielding consistent gains in sample efficiency, generalization to out-of-distribution problems, and training stability. Notably, these improvements hold across both synthetic tasks and real-world language reasoning benchmarks, highlighting the robustness and practical effectiveness of the proposed framework.
@@ 2.2 Reinforcement Learning for LLMs
Reinforcement learning from human feedback [24]—as popularized by InstructGPT—uses on-policy methods like PPO to fine-tune language models against a learned reward signal, but doing so requires costly fresh sampling, a separate value network and careful KL-penalty tuning, and remains sample-inefficient when rewards are asparse. [10, 25, 22] GRPO simplifies PPO by ditching the critic and computing advantages via within-prompt reward normalization across multiple candidates, yet it still depends on large volumes of model-generated outputs and can wander from the pre-trained behavior in low-reward regimes. These limitations—computational overhead, sample inefficiency and instability under sparse feedback—have spurred interest in leaner or hybrid approaches that retain the simplicity of policy gradients while injecting external guidance to stabilize and steer training.[8, 22]
@@ 2.3 Unifying SFT and RL
Unlike previous approaches that separate RL and SFT into distinct stages, SuperRL treats them as complementary processes governed by the data itself. It selectively applies offline supervision only when online rollouts lack useful reward signals—achieving more stable optimization and better generalization, particularly in tasks where sparse rewards make RL alone unreliable.
@@ 3 Methodology
3.1 Reward-Guided Loss Selection
We introduce a simple yet effective principle: when any rollout trajectory yields a nonzero reward, we apply a standard policy-gradient update to reinforce beneficial behaviors. However, if all trajectories return zero reward—signifying complete failure—the model instead performs a supervised update using offline demonstrations. This fallback mechanism ensures continuous learning even in the absence of meaningful reward signals.
Our approach treats the presence or absence of reward as a reliable indicator of learning potential. Rather than relying on hand-tuned heuristics or reward-density estimators, the rollout outcomes themselves determine whether reinforcement learning (RL) or supervised fine-tuning (SFT) is appropriate. This design is lightweight, hyperparameter-free, and operates at the per-instance level.
By selectively activating SFT only on challenging examples, the model allocates RL updates to solvable cases while directing supervision toward failure modes. This dynamic allocation improves training stability and accelerates convergence. On easier datasets, SFT is rarely triggered; under sparse-reward regimes, SFT becomes an automatic fail-safe. The result is a unified learning rule that is both simple and robust—consistently outperforming pure RL across diverse reasoning tasks.
SuperRL embodies this philosophy by integrating RL and SFT into a single reward-aware optimization loop that dynamically adjusts the training signal based on rollout feedback.
3.2 Reward-Gated Optimization in SuperRL
Notation.
For a task instance $x\!\in\!\mathcal{X}$ we draw $K$ trajectories $\tilde{y}_{1},\dots,\tilde{y}_{K}\!\sim\!\pi_{\theta}(\cdot\mid x)$ and obtain rewards $R(x,\tilde{y}_{k})$. A demonstration set $\mathcal{Y}_{\mathrm{SFT}}(x)=\{y^{*(1)},\dots,y^{*(M)}\}$ is available for imitation. Define the advantage and importance ratio
$A_{k}=R(x,\tilde{y}_{k})-b(x),\qquad r_{k}(\theta)=\frac{\pi_{\theta}(\tilde{y}_{k}\!\mid\!x)}{\pi_{\theta_{\mathrm{old}}}(\tilde{y}_{k}\!\mid\!x)}.$
Two primitive losses.
$\mathcal{L}_{\mathrm{SFT}}(\theta;x)$ $=-\frac{1}{M}\sum_{m=1}^{M}\log\pi_{\theta}\!\bigl{(}y^{*(m)}\mid x\bigr{)},$ (1)
$\mathcal{L}_{\mathrm{PG}}(\theta;x)$ $=-\frac{1}{K}\sum_{k=1}^{K}g\!\bigl{(}r_{k}(\theta),A_{k}\bigr{)},$ (2)
where $g(\cdot,\cdot)$ is any monotone policy-gradient surrogate (e.g. PPO, TRPO, GRPO).
Reward-based soft gate.
We define a soft gate that checks whether any rollout receives nonzero reward:
$c(x)=\mathbf{1}\!\bigl{(}\max_{k}R(x,\tilde{y}_{k})>0\bigr{)}\;\qquad\text{(triggered if any rollout succeeds)}$
Based on this gate, we define the unified objective:
$\boxed{\;\mathcal{L}_{SuperRL}(\theta;x)=(1-c(x))\,\mathcal{L}_{\mathrm{SFT}}(\theta;x)+c(x)\,\mathcal{L}_{\mathrm{PG}}(\theta;x)\;}$
If every rollout fails ($c(x)=0$), the model performs supervised imitation using demonstrations; otherwise, it reinforces successful behaviors via policy gradients. This hyperparameter-free switch tightly couples exploration and imitation, providing stable and sample-efficient learning that outperforms pure RL across both dense- and sparse-reward reasoning tasks.
3.3 Design Variants and Unified Hybrid Objective
The hard switch in Eq. (3.2) forces every instance to be handled by either supervised imitation or policy gradients. We explore two relaxations that keep the notation of Sec. 3.2—namely $A_{k},\;r_{k},\;g(\cdot,\cdot)$—while softening the coupling between the two learning signals.
Advantage-based soft gate.
We define an alternative binary gate that triggers RL only if at least one trajectory has a strictly positive advantage:
$c_{\!\mathcal{A}}(x)=\mathbf{1}\!\bigl{(}\max_{k}A_{k}>0\bigr{)}\qquad\text{(triggered if any rollout outperforms the baseline)}$
The resulting loss is:
$\boxed{\;\mathcal{L}_{\text{AdvGate}}(\theta;x)=(1-c_{\!\mathcal{A}}(x))\,\mathcal{L}_{\mathrm{SFT}}(\theta;x)+c_{\!\mathcal{A}}(x)\,\mathcal{L}_{\mathrm{PG}}(\theta;x)\;}$
Unlike the original SuperRL switch, which triggers RL updates whenever any rollout yields a nonzero raw reward, the advantage-based gate examines whether any trajectory yields a strictly positive advantage—that is, outperforms a learned baseline $b(x)$. This makes it more selective: dense but uninformative reward patterns (e.g., constant shaping terms) are often cancelled out by the baseline, resulting in $A_{k}=0$ and thus closing the gate. Even when raw rewards are nonzero, the update may fall back to SFT if no genuine improvement is detected. This design prevents futile RL updates on deceptively positive but sub-optimal trajectories and is particularly effective in sparse-reward settings with frequent shaping noise or misleading reward signals.
Uncertainty-weighted hybrid fusion.
A fully continuous alternative removes the binary gate and learns how much to trust each loss through two log-variance scalars $\sigma_{\mathrm{pg}}$ and $\sigma_{\mathrm{sft}}$:
$\boxed{\mathcal{L}_{\text{Hybrid}}(\theta)=e^{-2\sigma_{\mathrm{pg}}}\,\mathcal{L}_{\mathrm{PG}}(\theta)+e^{-2\sigma_{\mathrm{sft}}}\,\mathcal{L}_{\mathrm{SFT}}(\theta)+\sigma_{\mathrm{pg}}+\sigma_{\mathrm{sft}}}$
where $\mathcal{L}_{\mathrm{PG}}$ may be instantiated by PPO, TRPO, or GRPO. Demonstration and on-policy batches are mixed within every update; the learned variances down-weight noisy gradients yet the additive $\sigma$-terms stop either branch from collapsing to zero. During training, we interleave high-quality demonstrations with roll-outs in each mini-batch. The two uncertainty parameters adaptively balance signal quality without any explicit switching rule: noisy or unstable gradients are automatically suppressed by the factors $e^{-2\sigma}$, while the additive $\sigma$ penalties prevent the model from becoming overconfident.
Equation (1) preserves the hyperparameter-free spirit of the original gate while being more selective; Eq. (2) offers a differentiable, learnable fusion at the cost of two extra parameters.